Selected account balances for the year ended December 31 are provided below for Superior Company:

  
Selling and administrative salaries $ 105,000
Insurance, factory 6,200
Utilities, factory 45,500
Purchases of raw materials ?
Indirect labour 62,000
Direct labour 95,900
Advertising expense 77,200
Cleaning supplies, factory 7,400
Sales commissions 50,500
Rent, factory building 116,000
Maintenance, factory 30,000


Inventory balances at the beginning and end of the year were as follows:

Beginning of
the Year
End of
the Year
Raw materials $ 50,000 $ 11,500
Work in process ? 32,100
Finished goods 46,000 86,000

   

The total manufacturing costs for the year were $703,000, and the cost of goods sold totalled $664,000.

Required:
1.a.
Prepare a schedule of cost of goods manufactured.



1.b. Prepare the cost of goods sold section of the company's income statement for the year.



2. Assume that the dollar amounts given above are for the equivalent of 40,000 units produced during the year. Compute the average cost per unit for direct materials used and the average cost per unit for rent on the factory building.(Round your answers to 2 decimal places.)
  



3. Assume that in the following year the company expects to produce 50,000 units. What average cost per unit and total cost would you expect to be incurred for direct materials? For rent on the factory building? (Assume that direct materials are a variable cost and that rent is a fixed cost.) (Round your "Average Cost'' answers to 2 decimal places.)



4. This part of the question is not part of your Connect assignment.

Next Visit question map

Question1of5Total1 of 5

Prev
Answer #1

SOLUTION:

1.a. Prepare a schedule of cost of goods manufactured.

Schedule of Cost of Goods Manufactured
Particulars Amount Amount
Direct Materials:
Raw Materials Inventory Beginning (given) 50000
Add: Purchases of Raw Materials (Balancing Figure) 301500
Raw Materials available for use 351500
Less: Raw Materials Inventory Ending (given) 11500
Raw Materials used in Production (Balancing Figure) 340000
Direct labour (given) 95900
Manufacturing Overheas:
Indirect labour 62000
Insurance Factory 6200
Utilities Faactory 45500
Cleaning supplies factory 7400
Rent factory building 116000
Maintenance factory 30000
Total overhead cost ( all given) 267100
Total Manufacturing Cost (given) 703000
Add: WIP Inventory beginning (Balancing figure) 33100
736100
Less: WIP Inventory Ending (given) 32100
Cost of Goods Manufactured (From 1b below) 704000

1.b. Prepare the cost of goods sold section of the company's income statement for the year.

Schedule for Cost of Goods sold
Particulars Amount
Finished Goods Inventory Beginning (Given) 46000
Add: Cost of Goods Manufactured (Balancing Figure) 704000
Goods Available for Sale 750000
Less: Finished Goods Inventory Ending (Given) 86000
Cost of Goods Sold (Given) 664000

2. Assume that the dollar amounts given above are for the equivalent of 40,000 units produced during the year.

If 40000 units are produced,

then

average cost per direct materials = 340000/40000 = 8.5

average cost per rent factory building = 116000/40000 = 2.9

Hence,

Average Cost
Direct materials $ 8.5 per unit
Rent Factory Building $ 2.9 per unit

3. Assume that in the following year the company expects to produce 50,000 units. What average cost per unit and total cost would you expect to be incurred for direct materials? For rent on the factory building?

Since material cost is variable, it will change as per the number of units produced. However, rent, factory cost will remain the same i.e 116000

So average cost per direct material for 50000 units is 8.5 and average cost per rent, factory building is 116000/50000 = 2.32 per unit

Unit Cost Total Cost
Direct Materials $ 8.5 per unit $ 425000 (50000*8.5)
Rent Factory Building $ 2.32 per unit $ 116000 (50000*2.32)